OpenStudy (notamathgenius):

The ages of​ Edna, Ellie, and Elsa are consecutive integers. The sum of their ages is 114. What are their​ ages?

OpenStudy (benji):

Ah I love consecutive problems, so we have three this time, thinking about what we did last time what do you think it will look like if you start the equation as \[114=\]

satellite73 (satellite73):

i have an idea to heck with algebra, just guess

OpenStudy (benji):

haha

OpenStudy (notamathgenius):

I cant get @satellite73 this is for college :(

satellite73 (satellite73):

\[32+33+34=98\] too low, up it a bit

OpenStudy (benji):

use the equation \[144=x+x+1+x+2\]

OpenStudy (notamathgenius):

144= 3x+1+2

satellite73 (satellite73):

if you are in college, then you should be able to "think" i.e suss out a problem without being a slave to some method you will learn something as you do it

OpenStudy (benji):

Yep now solve for x and thats the youngest

satellite73 (satellite73):

114 right? not 144

OpenStudy (benji):

ya mb

OpenStudy (notamathgenius):

114******

satellite73 (satellite73):

which is why i guessed \[32+33+34\] but it didn't work, so try \[34+35+36\] and see if that one does

OpenStudy (notamathgenius):

its 105

OpenStudy (benji):

114-3 = 111/3=?

OpenStudy (notamathgenius):

i have no clue im mad now

OpenStudy (benji):

Alright let's take a step back and go to the original equation again\[114=3x+3\]

OpenStudy (notamathgenius):

ok so its 37

OpenStudy (benji):

yes that's the youngest, how old are the other two?

OpenStudy (notamathgenius):

38 39

OpenStudy (benji):

yep

OpenStudy (notamathgenius):

YAYAYAYAYAYA